Please complete all of #6 :) your result to that from the previous step). 6. Repeat the previous problem. This time, the stress state is oz = 10 ksi, oy =5. It has been determined that the critical stress in a part under a certain load is given by 0r = 10 ksi and oy = 20 ksi. Th
Answer #1

loksi 6 7 6y= - 2oksi Garg = 10 + (-20) _ - 5ksi R = Ex- Gavg) + Tay² = 5 (10-1-555 = 15ksi tman = R= 15KSI 6- Gaug th= -5 +1the factor of safety is more in distortion energy method than the maximum shear stress method.